OpenStudy (anonymous):

4√8x = 12√3x? correct? or not?

myininaya (myininaya):

\[4 \sqrt{8x}=4 \sqrt{8} \sqrt{x}= 4 \sqrt{4 \cdot 2} \sqrt{x} =4 \sqrt{4} \sqrt{2} \sqrt{x}=4 \sqrt{4} \sqrt{2} \sqrt{x}=4 (2) \sqrt{2x}=8 \sqrt{2x}\]

OpenStudy (anonymous):

wait. its 4√18 not 4√8

myininaya (myininaya):

\[4 \sqrt{18}=4 \sqrt{9} \sqrt{2}=4(3)\sqrt{2}=12 \sqrt{2}\]

myininaya (myininaya):

so the answer is still no

myininaya (myininaya):

you need 2 inside the square root not 3

OpenStudy (anonymous):

this is what i did.: 4√18x= 4√6*3x = 4*3√3x = 12√3x wher did i go wrong

OpenStudy (anonymous):

Not correct \[4\sqrt{8x}\] \[4\sqrt{4*2*x}\]\[4*2\sqrt{2x}\] \[8\sqrt{2x}\]

myininaya (myininaya):

yeah we already did that for that one but then she said it was 18 not 8 inside the square root